OpenStudy (anonymous):

1/3 - 3/10 = please help break down answer

OpenStudy (anonymous):

1/3 = 10/30 3/10= 9/30 therefore 1/30 would be the answer

OpenStudy (anonymous):

how did you get your answer?

OpenStudy (anonymous):

how did you get 10/30?

OpenStudy (anonymous):

You Have To Find The Least Common Denominator. In This Case, It's 30. 3x10 = 30 . 1 x 10 = 10 . 10/30 - 9/10 . Do You Get It ?

OpenStudy (anonymous):

* 9/30 .

OpenStudy (anonymous):

mulitply demoninator from both fractions?

OpenStudy (anonymous):

& The Numerator .

OpenStudy (anonymous):

well 1/3 you need a common denominator. and 30 is the common between the two (multiply both bottom numbers 3* 10) so you need 30 for both. to get 30 for the first one multiply both numbers on top and bottom by 10. for the second part multiply both numbers by 3 so that 3*3 is 9 and 3*10 =30 so what you are left with is 10/30 - 9/30

OpenStudy (anonymous):

so lost sorry. how did you get 10/30 - 9/30 which is the answer?

OpenStudy (anonymous):

To Subtract 1/3 - 3/10 . You Need A Least Common Denominator . 30 Is The Least Common Denominator. To Get The Denominator To 30 From 1/3, You Have To Multiply The Fraction By 10. So 1*10=10, 3*10=30 . So You Get 10/30 . Now You Do The Same To The Other Side . You Should Get 1/30 As The Answer Which Is Equal To 1/3 .

OpenStudy (anonymous):

the common denominator would be 30, right? so multiply 1/3*10/10 which gives u 10/30. multiplu 3/10 by 3/3 to get 9/30. so 10/30-9/30=1/30

OpenStudy (anonymous):

got it thank you so much